OpenStudy (anonymous):

Write the equation of the line with a slope of -1/3 and passing through the point (6, -4)

OpenStudy (anonymous):

Sure..... The equation of the straight line with slope "m" and passing through the point (x=0,y=b) is given by y=mx+b

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Here, we have m=-1/3, therefore, y=(-x/3)+b Now, the above line must pass through the point (x=6, y=-4)...substituting x and y in the linear equation we get -4=(-6/3)+b ---> -4=-2+b ---> b=-4+2 --> b=-2 Therefore, y=(-x/3)-2

OpenStudy (anonymous):

Write the equation of the line passing through the points (0, -4) and (-2, 2).

OpenStudy (anonymous):

dont u start by finding the slope? if u can please help me do this one :)

OpenStudy (anonymous):

The linear equation is given by y=mx+b Here, we are looking for the value of "m" and "b"....since you have two unknowns, you need two equations to solve for both unknowns. The linear equation must satisfy the fact that: when x=0, we have y=-4 so we get -4=m(0)+b ---> b=-4 Also, the line must pass through the point (-2, 2), i.e., when x=-2, we must have y=2, so 2=m(-2)+(-4) ---> m(-2)=6 ---> m=-3 therefore, y=-3x-4

OpenStudy (anonymous):

did u find the slope first ?

OpenStudy (anonymous):

Write the equation of the line passing through the points (-6,1) and (-4,2)

OpenStudy (anonymous):

I found "b" y-axis intersection point

OpenStudy (anonymous):

kay can u help me wwith one or two more please

OpenStudy (anonymous):

I did that because I have a point that pass through x=0

OpenStudy (anonymous):

Write the equation of the line passing through the points (-6,1) and (-4,2)

OpenStudy (anonymous):

every time you has a line that passes though a point with (0,y)...use it to find "b"

OpenStudy (anonymous):

For your last question, you need to find the slope first

OpenStudy (anonymous):

well can you help me please

OpenStudy (anonymous):

slope=m=(y2-y1)/(x2-x1)=(2-1)/(-4-(-6))=1/(-4+6)=1/2=0.5 So, we have now y=0.5x+b Now take any point one the line, say (-6,1) and plug it in the above equation, you get 1=0.5(-6)+b --> 1=-3+b --> b=4 Therefore, y=0.5x+4

OpenStudy (anonymous):

Write the equation of the line with an undefined slope, passing through the point (2, 5).

OpenStudy (anonymous):

A line with an undefined slope passing through a point (a,b) means a vertical line with equation x=a Here you have x=2

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

x=2 is the answer

OpenStudy (anonymous):

one last question :)

OpenStudy (anonymous):

okaaay...

OpenStudy (anonymous):